One plane ate 520 feet is ascending at a rate of 40 feet per minute, while another plane at 3800 feet is descending at a rate of 120 feet per minute. How long will it take the two planes to be at the same altitude?
Answer by mananth(16946)

One plane ate 520 feet is ascending at a rate of 40 feet per minute, while another plane at 3800 feet is descending at a rate of 120 feet per minute. How long will it take the two planes to be at the same altitude?
..
Plane A 520 feet vertical rise 40 feet/min
plane B 3800 feet vertical dip 120 feet per minute.
..
This means that the vertically the speeds are 40 feet / minute & 120 feet /minute travelling towards each other.
so Add up the speeds
..
120 + 40 = 160 feet per minute.
the distance between them = 3800 -520 =3280 feet
..
t= d/r
3280/160
t = 20.5 minutes. they will be at the same altitude.
...
m.ananth@hotmail.ca